# FIT-GNN: Faster Inference Time for GNNs Using Coarsening

### v2

@@ 4.1 Node-level Task
We use Algorithm 1 with $L$ number of layers to create a standard GNN model for node-level tasks. Training the model on $G^{\prime}$ is similar to the Huang et al. (2021) training algorithm. We design Algorithm 2 to train the model on $\mathcal{G}_{s}$.
For the node classification task, $\arg\max$ is used for creating labels for $G^{\prime}$ and CrossEntropy as a loss function. For the node regression task, we do not create any coarsened graph. Mean Absolute Error (MAE) is used as the loss function.
@@ 4.2 Graph-level Task
Unlike node-level tasks, the number of dataset samples is not decreasing. For a given graph $G\in\mathcal{D}$, we create $G^{\prime}$ and $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$. Each subgraph contains overlapping nodes when we use Extra Nodes or Cluster Nodes. Therefore, a boolean array $mask_{i}$ is created for each subgraph $G_{i}$ such that $True$ is set for nodes that actually belong to the subgraph $G_{i}$ (not as an Extra Nodes or Cluster Nodes); otherwise, $False$.
Unlike node-level tasks, we cannot train and infer on $G^{\prime}$ and $\mathcal{G}_{s}$ using the same model. We use Algorithm 3 to create a model to train on $G^{\prime}$. We design Algorithm 4 to train the model on $\mathcal{G}_{s}$. We use the CrossEntropy loss function for the graph classification task and the Mean Absolute Error(MAE) for the graph regression task. We use Algorithm 2 without the backpropagation for inference.
1:Number of Layers $L;A^{\prime};D^{\prime};X_{0}=X^{\prime}$
2:for $i=1$ to $L$ do
@@
4:end for
5:$\bar{X}=\textup{MaxPooling}(X_{L})$
6:$Z=\bar{X}\mathcal{W}_{L}$
7:return $Z$
Algorithm 4 Graph-Model-$\mathcal{G}_{s}(\mathcal{G}_{s},L,mask)$
1:$\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$; Number of Layers $L$; $mask=\{mask_{1},mask_{2},\dots,mask_{k}\}$;
2:$X\leftarrow[]$
@@
5:         $X_{j}=\sigma(\tilde{D_{i}}^{-\frac{1}{2}}\tilde{A_{i}}\tilde{D_{i}}^{-\frac{1}{2}}X_{j-1}\mathcal{W}_{j-1})$
@@
Train Test
@@
Coarsening $kd^{2}+k^{2}d$ $nd^{2}+n^{2}d$
FIT-GNN $\begin{matrix}kd^{2}+k^{2}d\\ +\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]\end{matrix}$ $\begin{matrix}\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]\end{matrix}$
Table 1: This table shows the training and inference time complexity of FIT-GNN as compared to classical and coarsening approaches.
Train Test
@@
Coarsening $k^{2}+kd+d^{2}$ $n^{2}+nd+d^{2}$
FIT-GNN $\begin{matrix}k^{2}+kd+d^{2}\\ +\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]\end{matrix}$ $\begin{matrix}d^{2}\\ +\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]\end{matrix}$
Table 2: This table shows the training and inference space complexity of FIT-GNN compared to classical and coarsening approaches.
4.3 Time and Space Complexity
The dimensions of the matrices in Equation 1 are $(n\times n)$ for $\tilde{A}$, $(n\times d)$ for $X_{l}$ and $(d\times d)$ for $\mathcal{W}_{l}$. Time complexity for one layer of GNN computation is $\mathcal{O}(n^{2}d+nd^{2})$. If we take $L$ layers, then the total time is $\mathcal{O}(Ln^{2}d+Lnd^{2})$. The space complexity is $\mathcal{O}(n^{2}+Lnd+Ld^{2})$. When we compute on a sparse graph, the time complexity is $\mathcal{O}(m+Lnd^{2})$ and space complexity is $\mathcal{O}(m+Lnd+Ld^{2})$.
Huang et al. (2021) improved the time and space complexity (mentioned in Table 1 and Table 2) for training the network by reducing the number of nodes from $n$ to $k$. However, the inference time and space complexity remain the same.
Let us compare three different models. One is the classical model, where no coarsening or partitioning is done; the second is the Huang et al. (2021) approach; third is our approach where from the graph $G$ we create $G^{\prime}$ and $\mathcal{G}_{s}$.
The inference time complexity of our model is $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$, where $\bar{n}_{i}=n_{i}+\phi_{\textup{max}}$ where $\phi_{\textup{max}}=\max_{i}(\mathcal{E}_{G_{i}})$ for Extra Nodes or $\phi_{\textup{max}}=\max_{i}(\mathcal{C}_{G_{i}})$ for Cluster Nodes is the maximum number of additional nodes added per subgraph. Table 1 and Table 2 show the comparison between different approaches.
Lemma 2.
The inference time complexity is at most $\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\leq n^{2}d+nd^{2}$ for $r\leq\frac{d-2}{d+\phi_{\textup{max}}}$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}$
@@ 4.3 Time and Space Complexity
$\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\\ =\sum_{i=1}^{k}[n_{i}^{2}d+\phi_{\textup{max}}^{2}d+2n_{i}\phi_{\textup{max}}d+n_{i}d^{2}+\phi_{\textup{max}}d^{2}]\\ =k\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+k\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d\\ =nr\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+nr\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d\\ (\textup{where }k=nr)\\ =nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d$
Now we know, $r\leq\frac{d-2}{d+\phi_{\textup{max}}}\Rightarrow r\phi_{\textup{max}}+2+rd\leq d$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}\Rightarrow n\phi_{\textup{max}}d\leq n^{2}-\sum_{i=1}^{k}n_{i}^{2}$.
Using this, we can say,
$\begin{split}&nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d\\ \leq&nd^{2}+(n^{2}-\sum_{i=1}^{k}n_{i}^{2})d+\sum_{i=1}^{k}n_{i}^{2}d\\ =&nd^{2}+n^{2}d\end{split}$
@@
For a given graph, if the above condition is satisfied, our approach has better inference time and space complexity than other approaches.
@@ 5 Experiments
Once we have constructed $G^{\prime}$ and $\mathcal{G}_{s}$, we train FIT-GNN in 4 different setups:
@@
$\bullet$ Gc-train-to-Gs-train: Train the GNN model on $G^{\prime}$, then use the learned weight as an initialization for training and final inference on $\mathcal{G}_{s}$.
@@
$\bullet$ Gc-train-to-Gs-infer: Train the GNN model on $G^{\prime}$, then infer on $\mathcal{G}_{s}$.
$\bullet$ Gs-train-to-Gs-infer: Train and infer only on $\mathcal{G}_{s}$.
@@
$\bullet$ Gc-train-to-Gc-infer: Unlike node-level tasks, we can use $G^{\prime}$ to infer for graph-level tasks. Therefore, in this setup, we train and infer on $G^{\prime}$.
@@
For tasks like node regression, we can only perform Gs-train-to-Gs-infer because a coarsened graph is not created.
5.1 Experimental Setup
Dataset descriptions are in Section A of the appendix, and Section C of the appendix covers the hyperparameters, device configuration, and key packages used for the experiments. Source Code present at https://github.com/Roy-Shubhajit/FIT-GNN.
5.1.1 Node regression
We use three real-world network datasets for the node regression task: Chameleon, Crocodile, and Squirrel Rozemberczki et al. (2019). The nodes are divided into train (30%), validation (20%), and test (50%) sets. We perform 20 runs and report the average and standard deviation of the smallest 10 normalized MAE, with normalization achieved by dividing the loss by the standard deviation of the target values.
5.1.2 Node classification
We conducted node classification experiments on six real-world network datasets: Cora McCallum et al. (2000), Citeseer Giles et al. (1998), Pubmed Sen et al. (2008), Coauthor Physics Shchur et al. (2018), DBLP Tang et al. (2008), and a subset of OGBN-Products Hu et al. (2020) extracted using Leiden algorithm Traag et al. (2019). We used three types of splits for the datasets: we applied the public split from Yang et al. (2016) for Cora, Citeseer, and PubMed. In the “few” split, we set aside labeled nodes per class for training and validation, while the rest form the test set. We ran 20 trials for each configuration and reported the average and standard deviation of the top 10 accuracies.
r Dataset
Chameleon Crocodile Squirrel
Classical - $0.844\pm 0.000$ $0.853\pm 0.000$ $0.809\pm 0.000$
FIT-GNN $0.1$ $\mathbf{0.460\pm 0.002}$ $\mathbf{0.362\pm 0.002}$ $\mathbf{0.648\pm 0.002}$
$0.3$ $0.499\pm 0.006$ $0.364\pm 0.001$ $0.672\pm 0.003$
Table 3: This table shows the normalized MAE loss (lower is better) for node regression tasks on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method.
Fixed Few
r Huang et al. (2021) FIT-GNN Huang et al. (2021) FIT-GNN
$0.1$ $0.772\pm 0.006$ $\mathbf{0.829\pm 0.005}$ $0.676\pm 0.051$ $\mathbf{0.695\pm 0.016}$
$0.3$ $\mathbf{0.817\pm 0.016}$ $0.815\pm 0.005$ $0.694\pm 0.045$ $\mathbf{0.719\pm 0.011}$
$0.5$ $\mathbf{0.827\pm 0.001}$ $0.807\pm 0.005$ $0.688\pm 0.045$ $\mathbf{0.729\pm 0.007}$
$0.7$ $\mathbf{0.824\pm 0.002}$ $0.800\pm 0.008$ $0.679\pm 0.046$ $\mathbf{0.706\pm 0.012}$
Table 4: Results on Cora Dataset. The metrics shown here include the average and standard deviation of the top 10 accuracies over 20 runs. We show Gc-train-to-Gs-train experimental setup result for fixed and Gc-train-to-Gs-infer experimental result for few. The highest accuracy for each coarsening ratio is shown in bold.
$\mathbf{r=0.1}$ $\mathbf{r=0.3}$
Dataset Split Huang et al. (2021) FIT-GNN Huang et al. (2021) FIT-GNN
Citeseer Few $0.583\pm 0.063$ $\mathbf{0.614\pm 0.021}$ $0.581\pm 0.052$ $\mathbf{0.624\pm 0.022}$
Fixed $\mathbf{0.711\pm 0.004}$ $0.677\pm 0.023$ $\mathbf{0.714\pm 0.003}$ $0.703\pm 0.007$
Pubmed Few $\mathbf{0.685\pm 0.052}$ $0.627\pm 0.116$ $\mathbf{0.687\pm 0.042}$ $0.676\pm 0.038$
Fixed $\mathbf{0.783\pm 0.005}$ $0.704\pm 0.022$ $\mathbf{0.784\pm 0.004}$ $0.677\pm 0.013$
DBLP Few $0.679\pm 0.056$ $\mathbf{0.683\pm 0.072}$ $0.648\pm 0.052$ $\mathbf{0.670\pm 0.034}$
Random $0.760\pm 0.021$ $\mathbf{0.789\pm 0.005}$ $0.745\pm 0.019$ $\mathbf{0.771\pm 0.007}$
Physics Coauthor Few $\mathbf{0.878\pm 0.036}$ $0.783\pm 0.028$ $\mathbf{0.908\pm 0.023}$ $0.877\pm 0.019$
Random $\mathbf{0.915\pm 0.014}$ $0.830\pm 0.048$ $\mathbf{0.934\pm 0.006}$ $0.910\pm 0.011$
OGBN-Products Random $0.296\pm 0.034$ $\mathbf{0.346\pm 0.014}$ $0.316\pm 0.016$ $\mathbf{0.406\pm 0.009}$
Table 5: Results for node classification tasks with accuracy as the metric (higher is better). We use the Cluster Nodes method to append additional nodes to subgraphs and Gc-train-to-Gs-train as experimental setup.
5.1.3 Graph Classification
Graph Classification has been conducted on the PROTEINS and AIDS dataset Morris et al. (2020); we randomly split the dataset into a $2:1:1$ ratio of train, validation, and test.
5.1.4 Graph Regression
@@
We perform graph regression on the QM9 Wu et al. (2017), and a subset of ZINC Gómez-Bombarelli et al. (2016) datasets, using the same splits as in graph classification. Gilmer et al. (2017) grouped the $19$ properties of each molecule in QM9 dataset into four broad categories. We predict one property from each of the four broad categories. The four specific properties are the following: the dipole moment ($\mu$), the gap between $\epsilon_{\mathrm{HOMO}}$ and $\epsilon_{\mathrm{LUMO}}$, zero-point vibration energy ($\mathrm{ZPVE}$), and at $298.15\mathrm{K}$, the atomization energy ($U^{\mathrm{ATOM}}$) of each molecule.
5.2 Result and Analysis

### v3

@@ 4.1 Node-level Task
We use Algorithm 4 with $L$ number of layers to create a standard GNN model $M$ for node-level tasks. While we use Algorithm 3 to train a model on the coarsened graph $G^{\prime}$, we propose Algorithm 1 to train a model on the set of subgraphs $\mathcal{G}_{s}$. Algorithm 3 utilizes the partition matrix $P$ generated by the coarsening algorithm to construct and train on the coarsened graph $G^{\prime}$ and the coarsened label information $Y^{\prime}$. On the other hand, our proposed Algorithm 1 uses the same partition matrix $P$ to construct and train on subgraphs $\mathcal{G}_{s}$ using original label information $Y$. For the node classification task, $\arg\max$ is used for creating labels for $G^{\prime}$ and CrossEntropy as a loss function. For the node regression task, we do not create $G^{\prime}$. Mean Absolute Error (MAE) is used as the loss function.
@@ 4.2 Graph-level Task
For a given graph $G\in\mathcal{D}$, we create $G^{\prime}$ and $\mathcal{G}_{s}=\{G_{1},G_{2},$ $\dots,G_{k}\}$. Since $G^{\prime}$ is a single graph and $\mathcal{G}_{s}$ is a set of subgraphs, we use Algorithm 5 to create a model to train on $G^{\prime}$ and Algorithm 2 to create a model to train on $\mathcal{G}_{s}$. We use the CrossEntropy loss function for the graph classification task and the Mean Absolute Error(MAE) for the graph regression task.
Algorithm 2 Graph-Model-$\mathcal{G}_{s}(\mathcal{G}_{s},L)$
@@
[…]
@@
[…]
@@
4.3 Time and Space Complexity
(a) Time Complexity
Train Inference (Full Graph)
@@ 4.3 Time and Space Complexity
SGGC $kd^{2}+k^{2}d$ $nd^{2}+n^{2}d$
FIT-GNN $\begin{matrix}kd^{2}+k^{2}d\\ +\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]\end{matrix}$ $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$
(b) Space Complexity
Train Inference
@@
SGGC $k^{2}+kd+d^{2}$ $n^{2}+nd+d^{2}$
FIT-GNN $\begin{matrix}k^{2}+kd+d^{2}\\ +\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]\end{matrix}$ $d^{2}+\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]$
Table 1: Training and inference time and space complexity of FIT-GNN compared to classical and SGGC approaches.
The dimensions of the matrices in Equation 1 are $(n\times n)$ for $\tilde{A}$, $(n\times d)$ for $X^{(l)}$ and $(d\times d)$ for $\mathcal{W}^{(l)}$. Time complexity for one layer of GNN computation is $\mathcal{O}(n^{2}d+nd^{2})$. If we take $L$ layers, then the total time is $\mathcal{O}(Ln^{2}d+Lnd^{2})$. The space complexity is $\mathcal{O}(n^{2}+Lnd+Ld^{2})$. When we compute on a sparse graph, the time complexity is $\mathcal{O}(m+Lnd^{2})$ and space complexity is $\mathcal{O}(m+Lnd+Ld^{2})$. The complexity is based on GCN. For other architectures like GAT [19], the complexity changes.
SGGC improved the time and space complexity (Table 1) for training the network by reducing the number of nodes from $n$ to $k$. However, the inference time and space complexity remain the same.
Let us compare three different models. One is the classical model, where no coarsening or partitioning is done; the second is SGGC, where training is performed on a smaller graph $G^{\prime}$ and inference on $G$; the third is our approach, FIT-GNN, where both training and inference are performed on $\mathcal{G}_{s}$.
The inference time complexity of our model is $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$, where $\bar{n}_{i}=n_{i}+\phi_{\textup{max}}$ where $\phi_{\textup{max}}=\max_{i}(\mathcal{E}_{G_{i}})$ for Extra Nodes or $\phi_{\textup{max}}=\max_{i}(\mathcal{C}_{G_{i}})$ for Cluster Nodes is the maximum number of additional nodes added per subgraph. Table 1 compares different approaches.
Lemma 4.2.
The inference time complexity is at most $\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\leq n^{2}d+nd^{2}$ for $r\leq\frac{d-2}{d+\phi_{\textup{max}}}$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}.$
@@
$\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]=\sum_{i=1}^{k}[n_{i}^{2}d+\phi_{\textup{max}}^{2}d+2n_{i}\phi_{\textup{max}}d+n_{i}d^{2}+\phi_{\textup{max}}d^{2}]$
$=k\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+k\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d$
$=nr\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+nr\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d\quad(\textup{where }k=nr)$
$=nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d$
Now we know, $r\leq\frac{d-2}{d+\phi_{\textup{max}}}\Rightarrow r\phi_{\textup{max}}+2+rd\leq d$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}\Rightarrow n\phi_{\textup{max}}d\leq n^{2}-\sum_{i=1}^{k}n_{i}^{2}$. Using this, we can say,
$nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d\leq nd^{2}+(n^{2}-\sum_{i=1}^{k}n_{i}^{2})d+\sum_{i=1}^{k}n_{i}^{2}d=nd^{2}+n^{2}d\\$
@@
For a given graph, when the above conditions are satisfied, our approach has better inference time and space complexity than other approaches. In Section C, we show that, for datasets with millions of nodes, the conditions outlined in Lemma 4.2 can indeed be satisfied. As the number of nodes increases, the gap in inference time between baseline methods and FIT-GNN widens. In table 9, we show the upper bound of $\phi_{\max}$ calculated as per the conditions mentioned in Lemma 4.2 for different datasets corresponding to different coarsening ratios. Negative upper bound is observed for some datasets, implying that our method won’t give better computational complexities in those scenarios.
@@ 5 Experiments
Let training on the coarsened graph $G^{\prime}$ be referred to as Gc-train, subgraph-level training as Gs-train, and subgraph-level inference process as Gs-infer. Once we have constructed $G^{\prime}$ and $\mathcal{G}_{s}$, we train FIT-GNN in 4 different setups:
@@
  • Gc-train-to-Gs-train: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then use the learned weight as an initialization for subgraph-level training and final inference on $\mathcal{G}_{s}$.
@@
  • Gc-train-to-Gs-infer: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then infer on $\mathcal{G}_{s}$ using the learned weights.
  • Gs-train-to-Gs-infer: Perform subgraph-level training and inference on $\mathcal{G}_{s}$.
@@
  • Gc-train-to-Gc-infer: Unlike node-level tasks, we can infer on $G^{\prime}$ (Gc-infer) itself for graph-level tasks after training on $G^{\prime}$. Therefore, in this setup for graph-level tasks only, we train and infer on $G^{\prime}$.
@@
All four setups can be applied to graph-level tasks. For the node classification task, we apply all setups except Gc-train-to-Gc-infer. For the node regression task, we only perform Gs-train-to-Gs-infer because a coarsened graph is not created for node regression datasets.
[…]
Tasks Datasets
Node Regression
(Train=30%, Val=20%, Test=50%)
Chameleon [20]
Crocodile [20]
Squirrel [20]
Node Classification
(Train=20*c, Val=30*c, Test=Rest)
Cora [21]
Citeseer [22]
Pubmed [23]
Coauthor Physics [24]
DBLP [25]
OGBN-Products [26]
Graph Regression
(Train=50%, Val=25%, Test=25%)
QM9 [27]
ZINC (Subset) [28]
Graph Classification
(Train=50%, Val=25%, Test=25%)
PROTEINS [29]
AIDS [29]
@@
Table 2: The table shows the datasets used and the splits in each dataset. ‘c’ represents the number of classes for the dataset. We applied the public split “fixed” from [30] for Cora, Citeseer, and PubMed. For the “random” split in node classification, we keep 2020 nodes per class for training, $30$ nodes for validation, and the rest as test nodes. [31] grouped the 1919 properties of each molecule in QM9 dataset into four broad categories. We predict one property from each broad category: the dipole moment ($\mu$), the gap between $\epsilon_{\mathrm{HOMO}}$ and $\epsilon_{\mathrm{LUMO}}$, zero-point vibration energy ($\mathrm{ZPVE}$), and at $298.15\mathrm{K}$, the atomization energy ($U^{\mathrm{ATOM}}$) of each molecule.
6 Result and Analysis
Methods Models
Reduction
Ratio (r)
Dataset
Chameleon Crocodile Squirel
Full GCN 1.0 $0.852\pm 0.001$ $0.854\pm 0.000$ $0.802\pm 0.000$
GAT 1.0 $0.846\pm 0.002$ $0.850\pm 0.000$ $0.805\pm 0.001$
SAGE 1.0 $0.848\pm 0.002$ $0.850\pm 0.000$ $0.804\pm 0.000$
GIN 1.0 $0.843\pm 0.000$ $0.852\pm 0.000$ $0.796\pm 0.000$
FIT-GNN GCN 0.1 $\mathbf{0.496\pm 0.005}$ $0.364\pm 0.001$ $0.663\pm 0.003$
GAT 0.1 $0.582\pm 0.029$ $0.395\pm 0.022$ $0.734\pm 0.019$
SAGE 0.1 $0.506\pm 0.006$ $\mathbf{0.309\pm 0.014}$ $\mathbf{0.589\pm 0.003}$
GIN 0.1 $0.769\pm 0.101$ $0.722\pm 0.098$ $0.795\pm 0.018$
Table 3: This table shows the normalized MAE loss (lower is better) for node regression task on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method. variation_neighborhoods coarsening algorithm is used.
Methods Models
Reduction
Ratio (r)
Dataset
Cora Citeseer Pubmed DBLP Physics
Full GCN 1.0 $0.821\pm 0.002$ $0.706\pm 0.002$ $0.768\pm 0.020$ $0.724\pm 0.002$ $\mathbf{0.933\pm 0.000}$
GAT 1.0 $0.809\pm 0.004$ $0.700\pm 0.003$ $0.740\pm 0.004$ $0.713\pm 0.004$ $0.909\pm 0.008$
SGGC GCN 0.3 $0.808\pm 0.003$ $0.700\pm 0.001$ $0.773\pm 0.002$ $0.737\pm 0.011$ $0.928\pm 0.003$
0.5 $0.808\pm 0.001$ $0.716\pm 0.002$ $\mathbf{0.793\pm 0.002}$ $0.733\pm 0.008$ $0.931\pm 0.004$
GAT 0.3 $0.686\pm 0.010$ $0.635\pm 0.021$ $0.773\pm 0.002$ $0.626\pm 0.028$ $0.870\pm 0.016$
0.5 $0.641\pm 0.026$ $0.635\pm 0.024$ $0.683\pm 0.024$ $0.604\pm 0.047$ $0.842\pm 0.018$
GCOND GCN 0.3 $0.806\pm 0.003$ $0.722\pm 0.001$ $0.779\pm 0.002$ $0.748\pm 0.001$ $0.876\pm 0.011$
0.5 $0.808\pm 0.004$ $\mathbf{0.730\pm 0.007}$ $0.773\pm 0.004$ $0.755\pm 0.002$ $0.804\pm 0.005$
GAT 0.3 $0.549\pm 0.080$ $0.628\pm 0.095$ $0.370\pm 0.079$ $0.481\pm 0.034$ OOM
0.5 $0.712\pm 0.059$ $0.501\pm 0.085$ $0.409\pm 0.044$ $0.464\pm 0.030$ OOM
BONSAI GCN 0.3 $0.722\pm 0.011$ $0.585\pm 0.003$ $0.538\pm 0.103$ $0.683\pm 0.039$ $0.798\pm 0.007$
0.5 $0.701\pm 0.012$ $0.666\pm 0.001$ $0.670\pm 0.084$ $0.702\pm 0.049$ $0.832\pm 0.016$
GAT 0.3 $0.680\pm 0.014$ $0.566\pm 0.003$ $0.540\pm 0.096$ $0.665\pm 0.027$ $0.791\pm 0.010$
0.5 $0.683\pm 0.006$ $0.579\pm 0.004$ $0.523\pm 0.133$ $0.681\pm 0.025$ $0.812\pm 0.016$
FIT-GNN GCN 0.3 $0.800\pm 0.003$ $0.679\pm 0.003$ $0.754\pm 0.001$ $\mathbf{0.786\pm 0.001}$ $0.932\pm 0.000$
0.5 $\mathbf{0.829\pm 0.002}$ $0.668\pm 0.003$ $0.761\pm 0.004$ $0.700\pm 0.001$ $0.926\pm 0.000$
GAT 0.3 $0.761\pm 0.004$ $0.655\pm 0.005$ $0.754\pm 0.002$ $0.771\pm 0.002$ $0.885\pm 0.004$
0.5 $0.792\pm 0.003$ $0.669\pm 0.004$ $0.760\pm 0.003$ $0.720\pm 0.002$ $0.885\pm 0.003$
Table 4: Results for node classification tasks with accuracy as the metric (higher is better). We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
6.1 Model Performance
Table 5: Results on OGBN-Products. FIT-GNN model uses a coarsening ratio $0.5$
Full $0.873\pm 0.000$
SGGC OOM
GCOND OOM
BONSAI OOM
FIT-GNN $\mathbf{0.894\pm 0.000}$
Table 3 shows the node regression error (mean and standard deviation of the lowest $10$ MAE) with coarsening ratios $0.1$. Results using different coarsening ratios are in Table 12. As the coarsening ratio increases, more subgraphs are created, and the size of each subgraph also reduces, and important neighboring node information gets lost. Hence, an increase in the coarsening ratio decreases performance.
Table 4 shows the results for node classification accuracy (mean and standard deviation of the top $10$ accuracies) with coarsening ratios $0.3$ and $0.5$ using the Cluster Nodes approach. Comparison with other coarsening ratios is in Table 13. Results show that the accuracy is comparable to the baselines. Table 5 shows results on the OGBN-Products dataset. Result on training on the full dataset from the OGBN-Products leaderboard [Link]. For training on the full dataset, $8\%$ for training, next top $2\%$ for validation, and the rest for testing. For our approach, we use $0.038\%$ training, $0.057\%$ validation, and the rest for testing.
Table 6 shows the results for the graph regression task on ZINC (Subset) and QM9 dataset. More results at Table 14. All the results show that the FIT-GNN model’s performance is better than the baselines. It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is high. It implies that finer subgraphs lose out on global information about the molecule, which is necessary for the prediction. Table 7 shows graph classification results. Training and inferring $G^{\prime}$ outperforms the baselines.
Methods Models
Reduction
Ratio (r)
Dataset
ZINC (Subset) QM9 (prop 0) QM9 (prop 4) QM9 (prop 6) QM9 (prop 13)
Full GCN 1.0 $0.743$ $0.855$ $1.012$ $1.102$ $1.073$
GAT 1.0 $0.736$ $0.919$ $1.015$ $1.104$ $1.066$
SAGE 1.0 $0.685$ $0.885$ $1.015$ $1.109$ $1.076$
GIN 1.0 $0.748$ $0.925$ $1.053$ $1.126$ $1.078$
FIT-GNN GCN 0.1 $0.625$ $0.713$ $1.223$ $\mathbf{0.447}$ $1.223$
GAT 0.1 $\mathbf{0.574}$ $\mathbf{0.705}$ $0.940$ $0.464$ $\mathbf{0.524}$
SAGE 0.1 $0.735$ $0.726$ $\mathbf{0.739}$ $0.526$ $0.905$
GIN 0.1 $0.644$ $0.726$ $0.746$ $0.494$ $0.856$
Table 6: Results on Graph Regression (lower is better); We use Extra node and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
Methods Models Graph per class Dataset
AIDS PROTEINS
DOSCOND GCN 1 $0.785$ $0.590$
10 $0.666$ $0.647$
50 $0.518$ $0.656$
GAT 1 $0.608$ $0.652$
10 $0.723$ $0.667$
50 $0.729$ $0.657$
KIDD GCN 1 $0.399$ $0.416$
10 $0.236$ $0.659$
50 $0.252$ $0.671$
GAT 1 $0.441$ $0.416$
10 $0.225$ $0.665$
50 $0.218$ $0.671$
SAGE 1 $0.653$ $0.416$
10 $0.361$ $0.650$
50 $0.283$ $0.673$
GIN 1 $0.759$ $0.412$
10 $0.637$ $0.655$
50 $0.634$ $0.674$
Methods Models
Reduction
Ratio (r)
Dataset
AIDS PROTEINS
Full GCN 1.0 $0.788$ $0.710$
GAT 1.0 $0.802$ $0.645$
SAGE 1.0 $0.770$ $0.613$
GIN 1.0 $0.800$ $0.774$
FIT-GNN GCN 0.1 $0.810$ $0.783$
0.3 $\mathbf{0.844}$ $\mathbf{0.826}$
0.5 $0.836$ $0.696$
0.7 $0.793$ $0.783$
GAT 0.1 $0.784$ $0.652$
0.3 $0.793$ $0.522$
0.5 $0.810$ $0.739$
0.7 $0.759$ $0.522$
SAGE 0.1 $0.793$ $0.696$
0.3 $0.828$ $0.696$
0.5 $0.793$ $0.696$
0.7 $0.819$ $\mathbf{0.826}$
GIN 0.1 $0.819$ $0.652$
0.3 $0.828$ $0.783$
0.5 $0.793$ $0.478$
0.7 $0.819$ $0.739$
Table 7: Results on Graph Classification (higher is better). We use the extra nodes method and Gc-train-to-Gc-infer as experimental setup. algebraic_JC coarsening algorithm is used.
6.2 Inference Time and Memory
As mentioned in Section 4.3, our inference time is less than the standard GNN if certain conditions are met. In Table 8(a) and 8(b), we club all baseline models into one, since all baselines are inferred on the whole graph. We empirically show the reduction of time and space by our FIT-GNN model during inference. For recording the inference time, we use the Python time package to calculate the difference in time before and after the inference step.